1) phosphorus pentachloride is produced by the reaction between phosphine and chlorine gas. What mass of chlorine gas would be required to produce 0.015kg of phosphorus pentachloride
2) if 10g of phosphine is made to react with 15g of chlorine gas. Calculate the percentage yield if 5.5g of phosphorus pentachloride was experimentally produced
1
Expert's answer
2019-04-01T07:48:43-0400

1)Reaction equation:

"PH_3+5Cl_2\to 2PCl_5 + 3H_2"

Chemical amount of phosphorus pentachloride :

"n=m\/M=0.015\/208.22 =7.2*10^{-5} mole"

To determine the chemical amount of chlorine gas we make up the proportion:


"x = 7.2*10^{-5} mole \\
5 mole = 2 mole"

The chemical amount of chlorine gas:

"x=5*7.2*10^{-5}\/2=18*10^{-5}mole"

Mass of chlorine gas:

"m(Cl_2)=n*M=18*10^{-5}*71=0.01278 g"


2)Find chemical amounts of phosphine and chlorine gas:


"n(PH_3)={\frac m M}={\frac {10} {34}}=0.294 mole"

"n(Cl_2)={\frac m M}={\frac {15} {71}}=0.211 mole"

Chemical amount of phosphorus pentachloride is determined by the chemical amount of chlorine, because it is in the minority:

"0.211 mole = x\\
5 mole = 2 mole"

The chemical amount of phosphorus pentachloride:


"x={\frac {0.211*2} {5}}=0.0844 mole"


Mass of phosphorus pentachloride


"m=n*M=0.0844*208.22=17.573g"


Percentage yield of phosphorus pentachloride:


"\eta ={\frac {5.5} {17.573}}*100\% =31.29\%"